# IAFormer: Interaction-Aware Transformer network for collider data analysis

### v1

@@ 2.3.2 Network structure and implementation
The IAFormer architecture comprises a data embedding block, attention-based layers, and a final MLP for classification as can be seen in Figure . Unlike original Transformer-based models, our approach does not require a class token to aggregate learned information across the layers. Instead, we replace it with a simple pooling operation applied to the final layer output.  [^1] The class token is particularly important in vision tasks, where an input image is divided into small patches, as it helps capture the global structure by aggregating information from Transformer layers. However, since our dataset consists of particle clouds, we replace the class token with average pooling [61]. Two input datasets are used, particle kinematic with dimension $(100,11)$ and interaction matrix with dimension $(100,100,6)$. Each dataset is processed through a dedicated embedding block. For the particle kinematics data, we employ an MLP consisting of three fully connected layers with 256, 128, and 32 neurons, respectively. Each layer is followed by a GELU activation function [62] and Root Mean Square Layer Normalization (RMSNorm) [63]. While different normalization techniques can be used, we found that RMSNorm enhances network convergence. For the pairwise interaction dataset, three two-dimensional convolution layers are employed with 256, 128, and 32 filters. The filter size is 1 such that the mapped features are not diluted by the convolution operation. Each convolution layer is followed by GELU activation and RMSNorm layers. The dimension of the embedded kinematic dataset is $(100,32)$ while the embedded particle interaction dataset has the dimension $(100,100,32)$.
The attention based layer has a distinct structure, processing two input datasets and outputting both datasets scaled by the computed attention. For each dataset, an RMSNorm layer is applied to normalize the input features. The trainable matrices, $W_{1}$ and $W_{2}$, are initialized using a fully connected layer for the particle kinematics dataset and a 2D convolutional layer for the pairwise interaction dataset. Both datasets are scaled using skip connections. For the particle kinematics dataset $X_{i}$, an MLP is added after the attention heads. This MLP consists of an RMSNorm layer followed by a fully connected layer with Sigmoid-Weighted Linear Units (SiLU) activation [64]. The MLP size is set to four times the number of embedded features. For the top tagging and quark-gluon tagging tasks, we use 12 attention layers, each containing 16 attention heads. The output of the final attention layer undergoes an average pooling before the final MLP. The final MLP layer consists of a fully connected layer with 100 neurons and an output layer with one neuron and a sigmoid activation function.
@@ 3.1 Input variables and data structure
In this analysis, we consider two input datasets, one with the particle momentum features and another one of particle pair interaction features. For the top tagging, up to $200$ constituent particles (hadrons) are retained for each jet, with the 4-momenta $(px,py,pz,E)$ of each particle. From this dataset, we construct the first input dataset with up to $100$ $p_{T}$ ordered jet constituents with eleven features for each particle as:
@@
$\Delta\eta=\eta-\eta_{\text{jet}}$ : pseudorapidity difference
$\Delta\phi=\phi-\phi_{\text{jet}}$ : azimuthal angle difference
@@
$\log\left(\frac{p_{T}}{p_{T_{\text{jet}}}}\right)$ : normalized $p_{T}$ (GeV)
$\log\left(\frac{E}{E_{\text{jet}}}\right)$ : normalized energy (GeV)
@@
[…]
Similar to [49], we consider the logarithm of the pairwise interaction variables. Integrating the energy and momentum variables in the interaction matrix is crucial as IAFormer uses the interaction matrix only for the attention score. The first two variables identify the most energetic particle pairs that most likely form the prongs of the Top jet.
3.2 Top tagging
The identification of jets originating from hadronically decaying top quarks, known as top tagging, plays a vital role in LHC new physics searches. To evaluate the performance of our proposed network, we employ the top tagging dataset. This dataset consists of jets generated at a center-of-mass energy of $\sqrt{s}=14$ TeV using Pythia8 [67], with fast detector simulation by Delphes [68]. The simulation does not incorporate effects from multiple parton interactions or pileup. Jet clustering is carried out using the Anti-kt algorithm with a cone radius of $R=0.8$, based on Delphes E-Flow objects. The dataset includes jets with transverse momentum in the range $p_{T}\in[550,650]$ GeV and pseudo-rapidity constrained to $|\eta|<2$. For top quark events, a valid jet must be located within $\Delta R=0.8$ of a hadronically decaying top quark, with all three decay products of the top also confined within $\Delta R=0.8$ from the jet axis. The primary background process considered is QCD dijet production. This dataset consists of one million $t\bar{t}$ events and an equal number of QCD dijet events. Following the standard data split, we allocate 1.2 million events for training, 400,000 for validation, and 400,000 for testing. The dataset is widely used in previous studies, allowing for direct performance comparisons of existing models.  [^2] It should be noted that there is an effective class imbalance near the top quark mass. The $t\bar{t}$ sample exhibits a pronounced peak around 170 GeV, whereas the mass of the QCD jets is concentrated near zero, resulting in tiny overlap between their distributions. This disparity makes it particularly challenging to discern fine-grained differences among high-performance neural networks. Therefore we use several measures to compare the network performance.
Table 1: Performance of different networks for Top jet classification. Results for networks without an asterisk are taken from their respective references. The plain Transformer was trained from scratch using the structure described in Appendix A.
ParticleNet[54] 0.940 0.9858 $397\pm 7$ $1615\pm 93$ 370K
PFN[53] -- 0.9819 $247\pm 3$ $888\pm 17$ 86.1K
rPCN[69] -- 0.984 $364\pm 9$ $1642\pm 93$-
Lorentz invariance based networks
PELICAN[35] 0.9426 0.987 -- $2250\pm 75$ 208K
LorentzNet[70] 0.942 0.9868 $498\pm 18$ $2195\pm 173$ 224K
L-GATr[71] 0.942 0.9870 $540\pm 20$ $2240\pm 70$
Attention based networks
ParT[49] 0.940 0.9858 $413\pm 6$ $1602\pm 81$ 2.14M
MIParT[50] 0.942 0.9868 $505\pm 8$ $2010\pm 97$ 720.9K
Mixer[21] 0.940 0.9859 $416\pm 5$ 86.03K
OmniLearn[72] 0.942 0.9872 $568\pm 9$ $2647\pm 192$ 1.6M
Plain Transformer 0.927 0.979 $362\pm 7$ $780\pm 73$ 1.7M
IAFormer 0.942 0.987 $510\pm 6$ $2012\pm 30$ 211K
Table 1 presents the classification performance of IAFormer for top jet classification. IAFormer is trained for 23 epochs, with a batch size of 256 with early stopping set for 5 epochs. The AdamW optimizer [73] with an initial learning rate of $5\times 10^{-4}$ and the Cosine annealing scheduler is used to adjust the learning rate during the training. Two input datasets are considered, particle kinematics, with dimensions $X=(100,11)$, with one-hundred maximum number of particles and 11 features per particle. A second dataset of the pairwise interaction $\cal{I}$, with dimension $(100,100,6)$, which encodes 6 features for each particle pair. Other networks are trained with the same hyperparameter setup, but trained for 20 epochs. After training, the network performance is evaluated using various metrics, including classification accuracy, Area Under the ROC Curve (AUC), and background rejection at signal efficiencies of 0.3 and 0.5.
IAFormer outperforms other attention-based Transformer networks, although it has an order of magnitude smaller parameter size of 211K than ParT. This reduction in network size is achieved by replacing the $Q$ and $K$ matrices with an interaction matrix, where attention scores are primarily based on pairwise particle interactions. Furthermore, the use of sparse attention enables the network to suppress attention scores of less relevant tokens, reducing the need for excessive model complexity, while efficiently distinguishing between Top and QCD jets.
Figure 2: Left plot shows the $\beta$ distribution across the twelve IAFormer layers for different seed numbers on the test dataset. Right plot illustrates the uncertainty in the network output when trained with different seed numbers, 42, 0, 7, 123, and 12345. The architectures of the Transformer $+I_{ij}$ and the plain Transformer. The uncertainty is reported in terms of background rejection at 0.3 signal efficiency.
This is reflected in the network training uncertainty in which IAFormer is much robust against numerical fluctuation. The right plot of figure 2 illustrates the fluctuation of the background rejection at 0.3 signal efficiency with the different seeds for the network parameter initialization. Furthermore, IAFormer exhibits robustness against the change of the seed number with a fluctuation range of $150$ while plain Transformer has a fluctuation range of $600$ and Transformer with interaction matrix has a fluctuation range of $550$.
The core component of the dynamic sparse attention is the inclusion of the learnable parameter $\beta$, which regulates the level of suppression of less relevant tokens. The left plot in Figure 2 illustrates the distribution of $\beta$ across all IAFormer layers. To better understand the role of $\beta$, we analyze three different random seeds. Interestingly, all distributions exhibit a similar pattern: $\beta$ values increase in the initial layers, reach a maximum value and then decrease in the later layers. Network classification accuracy improves for higher $\beta$; the blue line corresponds to the best classification accuracy of AUC=0.98678, with $\beta$ value very close to 0.6. It also starts at a lower value for earlier layers. This shows the earlier layer tries to build collective quantities stable against the fluctuations, and the rest of the information is abandoned in the later layers dynamically for successful training.
3.3 Quark gluon tagging
The quark jet and gluon jet are expected to show distinguishable features due to different fragmentation patterns, due to the color difference. Because BSMs tend to emit high-energy quarks rather than gluons, the quark-gluon tagging has been considered to be phenomenologically important. The public dataset [66] is generated using Pythia8 [67], where quark jets originate from $Z^{0}$ boson production plus quark processes, while gluon jets stem from $Z^{0}$ boson production plus gluon events. The dataset contains not only the four-momentum information but also particle id (PID), which can significantly enhance jet identification performance. The PID information is incorporated in a way that aligns with the experimental setup, categorizing particles into five broad types: electrons, muons, charged hadrons, neutral hadrons, and photons, along with their electric charge. These six PID attributes are incorporated as one hot vector, and they are combined with the seven kinematic variables, resulting in a total dimension of 13 input features per particle. These final state particles are clustered into jets using the anti-$k_{T}$ algorithm with a radius parameter of $R=0.4$. Only jets with transverse momentum $p_{T}$ in the range [500, 550] GeV and rapidity $|y|<2$ are included. The dataset consists of 2 million jets, evenly split between signal and background. Following the recommended partitioning, we allocate 1.6 million jets for training, 200,000 for validation, and 200,000 for testing for IAFormer and plain Transformer.
Table 2: Performance of different networks on the quark-gluon classification task. Results for networks without an asterisk are taken from their respective references. The plain Transformer was trained from scratch using the architecture described in Appendix A. The results for ParT and MIParT are shown in italic letters, because they include additional scaling factors for charged and neutral hadrons[49], which are not used in the other networks.
ParticleNet[54] 0.840 0.9116 $39.8\pm 0.2$ $98.6\pm 1.3$ 370K
PFN[53] -- 0.9052 $37.4\pm 0.7$- 86.1K
rPCN[69] -- 0.9081 $38.6\pm 0.5$- --
Lorentz invariance based networks
PELICAN[35] 0.8551 0.9252 ${52.3\pm 0.3}$ ${149.8\pm 2.4}$ 208K
LorentzNet[70] 0.844 0.9156 $42.4\pm 0.4$ $110.2\pm 1.3$ 224K
Attention based networks
ParT[49] 0.849 0.9203 $\mathit{47.9\pm 0.5}$ $\mathit{129.5\pm 0.9}$ 2.14M
MIParT[50] 0.851 0.9215 $\mathit{49.3\pm 0.4}$ $\mathit{133.9\pm 1.4}$ 720.9K
@@ 3.3 Quark gluon tagging
Plain Transformer 0.831 0.890 $32.6\pm 0.5$ $86.\pm 1.7$ 1.7M
IAFormer 0.844 0.917 $42\pm 0.3$ $101\pm 1.1$ 171K
[…]
@@
Figure 3: $\beta$ distribution for quark-gluon test data set for three different seed number.
@@
IAFormer architecture was optimized for best performance using 6 attention layers, with a total parameter count of 171k. The reduction in network size, compared to the top tagging case, is due to the similarity between quark and gluon jets. In this scenario, the network performance saturates after a certain number of layers, and adding more layers only introduces redundancy without improving accuracy. The optimization of the network size is based on the distribution of $\beta$ parameters for the attention layers. For IAFormer with 12 layers, $\beta$ distribution randomly fluctuates across layers with little improvement in the classification accuracy. By reducing the network size, the $\beta$ parameter shows a similar distribution as the Top tagging case, as shown in Figure 3.
@@
Finally, the obtained $\beta$ value of the last layer is larger than the top vs QCD classification of the previous subsections. Because $\beta$ appears negatively in the attentions, the larger $\beta$ likely indicates that the effective degree required for the quark-gluon classification is smaller than the top-QCD classification.

### v2

@@ 2.3.2 Network structure and implementation
The IAFormer architecture comprises a data embedding block, attention-based layers, and a final MLP for classification, as can be seen in Figure 1. Unlike original Transformer-based models, our approach does not require a class token to aggregate learned information across the layers. Instead, we replace it with a simple pooling operation applied to the final layer output [^1] The class token is particularly important in vision tasks, where an input image is divided into small patches, as it helps capture the global structure by aggregating information from Transformer layers. However, since our dataset consists of particle clouds, we replace the class token with average pooling [44]. . Two input datasets are used, particle kinematics with dimension $(100,11)$ and interaction matrix with dimension $(100,100,6)$. Each dataset is processed through a dedicated embedding block. For the particle kinematics data, we employ an MLP consisting of three fully connected layers with 256, 128, and 32 neurons, respectively. Each layer is followed by a GELU activation function [46] and Root Mean Square Layer Normalization (RMSNorm) [81]. While different normalization techniques can be used, we found that RMSNorm enhances network convergence. For the pairwise interaction dataset, three two-dimensional convolution layers are employed with 256, 128, and 32 filters. The filter size is 1 so that the mapped features are not diluted by the convolution operation. Each convolution layer is followed by GELU activation and RMSNorm layers. The dimension of the embedded kinematic dataset is $(100,32)$ while the embedded particle interaction dataset has the dimension $(100,100,32)$.
The attention-based layer has a distinct structure, processing two input datasets and outputting both datasets scaled by the computed attention. For each dataset, an RMSNorm layer is applied to normalize the input features. The trainable matrices, $W_{1}$ and $W_{2}$, are initialized using a fully connected layer for the particle kinematics dataset and a 2D convolutional layer for the pairwise interaction dataset. Both datasets are scaled using skip connections. For the particle kinematics dataset $X_{i}$, an MLP is added after the attention heads. This MLP consists of an RMSNorm layer followed by a fully connected layer with Sigmoid-Weighted Linear Units (SiLU) activation [29]. The MLP size is set to four times the number of embedded features. For the top tagging and quark-gluon tagging tasks, we use 12 attention layers, each containing 16 attention heads. The output of the final attention layer undergoes an average pooling before the final MLP. The final MLP layer consists of a fully connected layer with 100 neurons and an output layer with one neuron and a sigmoid activation function.
@@ 3.1 Input variables and data structure
In this analysis, we consider two input datasets, one of which is the particle momentum features and the other is the particle pair interaction features. For the top tagging, The original data set retains up to $200$ constituent 4-momenta $(px,py,pz,E)$ for each jet. For top jet tagging, we construct the input dataset with up to $100$ $p_{T}$ ordered jet constituents with eleven features as:
$\text{- }P_{4}=(p_{x},p_{y},p_{z},E)$ : particle 4-momentum (10)
$\text{- }\Delta\eta=\eta-\eta_{\text{jet}}$ : pseudorapidity difference
$\text{- }\Delta\phi=\phi-\phi_{\text{jet}}$ : azimuthal angle difference
$\text{- }\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}$ : angular distance from jet axis
$\text{- }\log(p_{T})$ : transverse momentum (GeV)
$\text{- }\log(E)$ : energy (GeV)
$\text{- }\log\left(\frac{p_{T}}{p_{T_{\text{jet}}}}\right)$ : normalized $p_{T}$ (GeV)
$\text{- }\log\left(\frac{E}{E_{\text{jet}}}\right)$ : normalized energy (GeV)
$\text{- }(p_{T_{a}}+p_{T_{b}})/p_{T_{j}}$ : sum of pair transverse momenta normalized by jet $p_{T}$ (11)
$\text{- }(E_{a}+E_{b})/E_{j}$ : sum of pair energies normalized by jet energy
$\text{- }\Delta=\sqrt{(\eta_{a}-\eta_{b})^{2}+(\phi_{a}-\phi_{b})^{2}}$ : angular distance between particles
$\text{- }k_{T}=\min(p_{T_{a}},p_{T_{b}})\cdot\Delta$ : transverse momentum scale
$\text{- }z=\min(p_{T_{a}},p_{T_{b}})/p_{T_{a}}+p_{T_{b}}$ : momentum sharing fraction
$\text{- }m^{2}=(E_{a}+E_{b})^{2}-\left|\mathbf{p}_{a}+\mathbf{p}_{b}\right|^{2}$ : invariant mass squared of the pair
Similar to [69], we consider the logarithm of the pairwise interaction variables. Integrating the energy and momentum variables in the interaction matrix is crucial, as IAFormer uses the interaction matrix only for the attention score. The first two variables identify the most energetic particle pairs that most likely form the prongs of the top jet.
Note that the interaction matrix is calculable from the particle kinematic dataset; however, in order to prioritize the inputs of the interaction matrix, one needs the knowledge of boost invariance and QCD, and a huge dataset is required for a plain Transformer to find the relevant information. The setup of IAFormer is therefore more efficient compared with that of Plain transformer.
3.2 Top tagging
The identification of jets originating from hadronically decaying top quarks, known as top tagging, plays a vital role in LHC new physics searches. To evaluate the performance of our proposed network, we employ the top tagging dataset. This dataset consists of jets generated at a center-of-mass energy of $\sqrt{s}=14$ TeV using Pythia8 [12], with fast detector simulation by Delphes [26]. The simulation does not incorporate effects from multiple parton interactions or pileup. Jet clustering is carried out using the Anti-$k_{T}$ algorithm with a radius parameter of $R=0.8$, based on Delphes E-Flow objects. The dataset includes jets with transverse momentum in the range $p_{T}\in[550,650]$ GeV and pseudo-rapidity constrained to $|\eta|<2$. For top quark events, a valid jet must be located within $\Delta R=0.8$ of a hadronically decaying top quark, with all three decay products of the top also confined within $\Delta R=0.8$ from the jet axis. The primary background process considered is QCD dijet production. This dataset consists of one million $t\bar{t}$ events and an equal number of QCD dijet events. Following the standard data split, we allocate 1.2 million events for training, 400,000 for validation, and 400,000 for testing. The dataset is widely used in previous studies, allowing for direct performance comparisons of existing models.  [^2] It should be noted that there is an effective class imbalance near the top quark mass. The $t\bar{t}$ sample exhibits a pronounced peak around 170 GeV, whereas the mass of the QCD jets is concentrated near zero, resulting in a tiny overlap between their distributions. This disparity makes it particularly challenging to discern fine-grained differences among high-performance neural networks. Therefore, we use several measures to compare the network performance.
Table 1: Performance of different networks for top jet classification. Results for networks without an asterisk are taken from their respective references. Uncertainty bands are computed from three independent runs. Plain Transformer was trained from scratch using the structure described in Appendix A.
ParticleNet[68] 0.940 0.9858 $397\pm 7$ $1615\pm 93$ 370K
PFN[57] - 0.9819 $247\pm 3$ $888\pm 17$ 86.1K
rPCN[71] - 0.984 $364\pm 9$ $1642\pm 93$ -
Lorentz invariance based networks
PELICAN[15] 0.9426 0.987 - $2250\pm 75$ 208K
LorentzNet[35] 0.942 0.9868 $498\pm 18$ $195\pm 173 22$4K
L-GATr[16] 0.942 0.9870 $540\pm 20$ $2240\pm 70$ 1.1M
Attention based networks
PCT[62] 0.940 0.9855 $392\pm 7$ $1533\pm 101$ 193.3K
ParT[69] 0.940 0.9858 $413\pm 6$ $1602\pm 81$ 2.14M
MIParT[79] 0.942 0.9868 $505\pm 8$ $2010\pm 97$ 720.9K
Mixer[43] 0.940 0.9859 $416\pm 5$ 86.03K
fine-tuned networks
ParticleNet-f.t.[69] 0.942 0.9866 $487\pm 9$ $1771\pm 80$ 370K
ParT-f.t.[69] 0.944 0.9877 $691\pm 15$ $2766\pm 130$ 2.14M
MIParT-f.t.[79] 0.944 0.9878 $640\pm 10$ $2789\pm 133$ 2.38M
L-GATr-f.t.[16] 0.9446 0.98793 $651\pm 11$ $2894\pm 84$ 1.1M
OmniLearn[63] 0.942 0.9872 $568\pm 9$ $2647\pm 192$ 1.6M
Plain Transformer 0.938 0.9837 $381\pm 7$ $1350\pm 70$ 2.1M
IAFormer 0.942 0.9870 $510\pm 6$ $2012\pm 30$ 211K
@@ 3.2 Top tagging
IAFormer offers performance comparable to other attention-based Transformer networks, while requiring an order of magnitude fewer parameters (211K) than ParT. This reduction in network size is achieved by replacing the $Q$ and $K$ matrices with an interaction matrix, where attention scores are primarily based on pairwise particle interactions. Furthermore, the use of sparse attention enables the network to suppress attention scores of less relevant tokens, reducing the need for excessive model complexity, while efficiently distinguishing between top and QCD jets.
3.2.1 The role of sparse attention
The core component of the dynamic sparse attention is the inclusion of the learnable parameter $\beta$, which regulates the level of suppression of less relevant tokens. The left plot in Figure 2 illustrates the distribution of $\beta$ across all IAFormer layers. To better understand the role of $\beta$, we analyze three different random seeds. Interestingly, all distributions exhibit a similar pattern, $\beta$ values increase in the initial layers, reach a maximum value, and then decrease in the later layers. Moreover, the network classification accuracy improves for higher $\beta$; the blue line corresponds to the best classification accuracy of AUC=0.98678, with $\beta$ value very close to 0.6. Also, it starts at a lower value for earlier layers. This shows that the earlier layer tries to build collective quantities stable against fluctuations, and the rest of the information is abandoned in the later layers dynamically for successful training.
Figure 2: The left plot shows the $\beta$ distribution across the twelve IAFormer layers for different seed numbers on the test dataset. The right plot illustrates the uncertainty in the network output when trained with different seed numbers, 42, 0, 7, 123, and 12345. The results of the architectures of the Transformer $\mathcal{I}_{ij}$(similar to ParT) and the plain Transformer are shown. The uncertainty is reported in terms of background rejection at 0.3 signal efficiency.
The use of sparse attention allows for a substantial decrease in the number of trainable parameters while maintaining the classification performance. This parameter reduction effectively constrains the optimization landscape, leading to smoother convergence dynamics and more stable training behavior. As a result, the model becomes less sensitive to random initialization and stochastic mini-batch fluctuations, which in turn manifests as reduced variance across independent training runs. The right plot of figure 2 illustrates the fluctuation of the background rejection at 0.3 signal efficiency with different seeds for the network parameter initialization. IAFormer exhibits robustness against the change of the seed number with a fluctuation range of $150$ while plain Transformer has a fluctuation range of $600$, and Transformer with interaction matrix (ParT) has a fluctuation range of $550$. To further illustrate the impact of sparse attention, we calculate the background rejection for IAFormer at $\beta=0$, which exhibits a fluctuation range of 350. In fact, IAFormer resistance to fluctuations arises from its low number of trainable parameters, the structure which achieves higher classification performance through dynamic sparse attention. In contrast, reducing the number of trainable parameters in the other networks typically results in a significant drop in classification performance.
To further evaluate network efficiency, we compute the number of Floating Point Operations (FLOPs) for network layers. This requires understanding the computational complexity of the core operations in both plain Transformer and IAFormer models. The primary contributors to the computational cost are the self-attention mechanism and the feed-forward layers. The self-attention operation in a Transformer has a complexity of $O(L^{2}\cdot d_{\text{model}})$, where L is the sequence length and $d_{\text{model}}$ is the model dimension. The number of FLOPs, for non-embedding parameters, is defined as $C_{\rm forwrad}=2N+2n_{\rm attn}d_{\rm model}n_{ctx}$, with $N$ is the number of non-embedding parameters in the model, $n_{\rm attn}$ is the number of the attention layer, $d_{\rm model}$ is the hidden dimension of the model and $n_{ctx}$ is the number of tokens. As a result, for a particle tokens length, the number of FLOPs for IAFormer is 38 million while for the plain transformer is 300 million. These results highlight the substantial reduction in computational cost achieved by IAFormer. In particular, the model offers an order-of-magnitude improvement in efficiency compared to the plain Transformer. In addition, the training time per batch of 256 events on an NVIDIA RTX A6000 is 11 seconds for $\beta=0$ and increases to 16 seconds for $\beta\neq 0$. Although the inclusion of $\beta$ leads to a longer training time, it results in superior classification performance compared to the $\beta=0$ case.
In addition to FLOPs, we measure peak tensor memory usage, which provides a hardware-independent estimate of memory requirements. For the used IAFormer configuration used in this study, the peak memory during training in FP32 precision is approximately 19.3 GB . The parameter memory is negligible, and the dominant contribution arises from intermediate pairwise interaction tensors scaling as $O\left(BN^{2}\right)$.
3.3 Quark gluon tagging
The quark jet and gluon jet are expected to show distinguishable features due to different fragmentation patterns arising from the color difference. The quark-gluon tagging is phenomenologically important because BSM particles tend to emit high-energy quarks rather than gluons. The public dataset [55] is generated using Pythia8 [12], where quark(gluon) jets originate from $Z^{0}$ boson plus quark(gluon) process. The dataset contains not only the four-momentum information but also the particle ID (PID) of jet constituents, which significantly enhances jet identification performance. The PID information is incorporated in a way that aligns with the experimental setup, namely, categorizing particles into five broad types: electrons, muons, charged hadrons, neutral hadrons, and photons, along with their electric charge. These six PID attributes are incorporated as one-hot vectors, and they are combined with the seven kinematic variables, resulting in a total dimension of 13 input features per particle.
The final state particles are clustered into jets using the anti-$k_{T}$ algorithm with a radius parameter of $R=0.4$. Only jets with transverse momentum $p_{T}$ in the range [500, 550] GeV and rapidity $|y|<2$ are included. The dataset consists of 2 million jets, evenly split between signal and background. Following the recommended partitioning, we allocate 1.6 million jets for training, 200,000 for validation, and 200,000 for testing for IAFormer and plain Transformer.
[…]
ParticleNet[68] 0.840 0.9116 $39.8\pm 0.2$ $98.6\pm 1.3$ 370K
PFN[57] - 0.9052 $37.4\pm 0.7$ - 86.1K
rPCN[71] - 0.9081 $38.6\pm 0.5$ - -
Lorentz invariance based networks
PELICAN[15] 0.8551 0.9252 $\bm{52.3\pm 0.3}$ $\bm{149.8\pm 2.4}$ 208K
LorentzNet[35] 0.844 0.9156 $42.4\pm 0.4$ $110.2\pm 1.3$ 224K
Attention based networks
ABCNet[61] 0.840 0.9126 $42.6\pm 0.4$ $118.4\pm 1.5$ 230K
PCT[62] 0.841 0.9140 $43.2\pm 0.7$ $118\pm 2.2$ 193.3K
@@ 3.3 Quark gluon tagging
fine-tuned networks
@@
Plain Transformer 0.839 0.9104 $39.4\pm 0.7$ $109.\pm 1.9$ 2.1M
IAFormer 0.844 0.9172 $42\pm 0.3$ $101\pm 1.1$ 171K
Table 2 presents the performance of IAFormer on the quark-gluon tagging task, compared to current state-of-the-art networks. Results for models without an asterisk are taken directly from their published works, while the plain Transformer results are obtained by training the model from scratch. It is important to note that the ParT and MIParT models outperform most other networks, primarily because they apply additional scaling factors to charged and neutral hadrons. Specifically, charged hadrons are weighted as $N(|\textsc{pid}|==211)+N(|\textsc{pid}|==2212)\times 0.5+N(|\textsc{pid}|==321)\times 0.2$, while neutral hadrons are weighted as $N(|\textsc{pid}==130|)+N(|\textsc{pid}==2112|)\times 0.5$, where 130, 211, 321, 2112, 2211 stands for $K_{L}$, $\pi^{+}$, $K^{+}$, $n$ and $p$ respectively. These heuristic scaling factors enhance classification accuracy by approximately $1\%$ over the baseline.
@@
Figure 3: $\beta$ distribution for quark-gluon test data set for three different seed numbers.
@@
For quark-gluon tagging, the number of layers of the IAFormer architecture is reduced to six with a total parameter count of 171k. In this scenario, the network performance saturates after a certain number of layers, and adding more layers only introduces redundancy without improving accuracy. The reduction in network size, compared to the top tagging case, may be understood as the result of the similarity between quark and gluon jets. Our choice of the number of layers is based on the distribution of $\beta$ parameters for the attention layers. For IAFormer with 12 layers, $\beta$ distribution randomly fluctuates across layers with little improvement in the classification accuracy. By reducing the network size by half, the $\beta$ parameter shows a similar distribution as the top tagging case, as shown in Figure 3. We did not optimize the network parameter further.
@@
Finally, the obtained $\beta$ value of the last layer is larger than the top vs QCD classification of the previous subsections. Because $\beta$ appears negatively in the attentions, the larger $\beta$ likely indicates that the effective degree of freedom required for the quark-gluon classification is smaller than the top-QCD classification.
3.4 JetClass Dataset
The JetClass dataset [69] is one of the largest and most comprehensive publicly available datasets designed for machine learning applications in jet physics. It was developed to provide a large scale, standardized benchmark for training and comparing modern deep learning architectures in high energy physics. The dataset is based on simulated proton-proton collision events at a center of mass energy of 14 TeV, generated with Pythia8 for event generation and processed through Delphes3 for fast detector simulation. This ensures a realistic description of both the parton shower and detector response, while maintaining computational efficiency suitable for large scale studies.
The dataset comprises a total of 100 million jets divided equally among ten distinct jet classes, each corresponding to a specific partonic origin or heavy resonance decay topology. These include light quark and gluon jets, heavy-flavor jets, as well as jets originating from hadronic decays of the $W$ and $Z$ bosons, the top quark and the Higgs boson decaying to bottom quarks. Each jet is represented as a collection of its reconstructed constituents, described by low-level particle features such as transverse momentum $p_{T}$, pseudorapidity $\eta$, azimuthal angle $\phi$, and energy $E$. In addition, high-level jet observables such as jet mass and substructure variables are also provided, allowing for studies that combine low-level and high-level representations within a unified framework.
Table 3: Performance of different networks on the JetClass dataset using only 10M events.
$h\!\to\!\bar{b}b$ $h\!\to\!\bar{c}c$ $h\!\to\!gg$ $h\!\to\!4q$ $h\!\to\!l\nu\bar{q}q$ $t\!\to\!b\bar{q}q$ $t\!\to\!bl\nu$ $W\!\to\!\bar{q}q^{\prime}$ $Z\!\to\!\bar{q}q$
$\rm Rej_{50\%}$ $\rm Rej_{50\%}$ $\rm Rej_{50\%}$ $\rm Rej_{50\%}$ $\rm Rej_{99\%}$ $\rm Rej_{50\%}$ $\rm Rej_{99.5\%}$ $\rm Rej_{50\%}$ $\rm Rej_{50\%}$
ParticleNet(10M) [69] 5848 2070 96 770 2350 5495 6803 307 253
ParT(10M) [69] 8734 3040 110 1274 3257 12579 8969 431 324
MIParT-L(10M) [79] 8000 3003 112 1281 3650 16529 9852 440 336
L-GATr(10M).[16] 9804 3883 120 1791 4255 24691 13333 506 373
IAFormer-L(10M) 8264 2968 114 1280 3425 14314 8731 423 345
A key feature of JetClass lies in its unprecedented size and diversity. With ten million examples per category, the dataset provides sufficient statistical power to train deep architectures with millions of parameters. Its inclusion of multiple jet categories ensures that models trained on it can capture a wide range of radiation patterns, color flow structures and multi-prong topologies arising from different SM processes at once. Thus, the task here differs slightly from the Top and quark–gluon datasets, which are trained to distinguish between two distinct classes. In this case, the model must classify among ten different categories simultaneously. Accordingly, we modified the output layer to contain ten neurons.
To evaluate the performance of IAFormer on the JetClass dataset, we train the network on 10 million events and compare its background rejection efficiency with that of other Transformer-based architectures. Motivated by the Chinchilla scaling law for large language models [47], we increase the size of IAFormer to 890k parameters to ensure that the model can effectively capture the rich information contained in the JetClass dataset. In addition, we used the same optimizer and learning rate schedule as in the top tagging study.Table 3 summarizes the background rejection efficiencies for all classes, comparing IAFormer-L with MIPart-L and ParT, when trained on a dataset of size 10M.
In addition, we measure peak tensor memory usage during forward and backward propagation in FP32 precision. For the JetClass IAFormer configuration used in this study, the model requires approximately 9.15 GB of peak GPU memory. The parameter memory itself is negligible ( 4 MB), and the dominant contribution arises from intermediate activation tensors within the sparse attention blocks. Memory scales approximately linearly with the number of edges in the graph, consistent with the expected $O\left(BN\right)$ complexity.